estion 7 of 16> The quadratic formula is used to solve for x in equations taking the frm of a quadratic cquation, ar + br + e-O quadratic formula: x Solve for x in the expression using the quadratic formula 2x2 + 29x-8. Use at least three significant figures in each answer. and x = 5 6 8 9
Answer #1

2x^2 + 29 x - 8.1 = 0

Now, for the value of x

D = sqrt(b^2- 4ac)

D = sqrt(29^2 + 4 * 2 * 8.1)

D = 30.0965

x = (29 + 30.0965)/4 , (29 - 30.0965)/4

x = 14.7741 , - 0.2741

the value of x are 14.7741 and - 0.2741